OpenStudy (kweb2017):

If a line crosses the y-axis at (0, -3) and has a slope of 3, what is the equation of the line?

OpenStudy (anonymous):

use this formula to find the equation\[y=mx+c\]where\[y=vertical~axis\]\[x=horizontal~axis\]\[m=slope\]\[c=y-intercept\]

OpenStudy (anonymous):

u have the value of x and the value of y and also the slope

OpenStudy (anonymous):

substitute the value inot the formula to find the equation

OpenStudy (anonymous):

into*

OpenStudy (anonymous):

\[y=-3\]\[x=0\]\[m=3\]\[c=?\]

OpenStudy (anonymous):

\[y=mx+\color{red}{c}\] substitute the value to find c

OpenStudy (anonymous):

\[-3=3(0)+c\]

OpenStudy (kweb2017):

so it would be 3x+y=3?

OpenStudy (anonymous):

how do u get that?

OpenStudy (kweb2017):

its one of my options on my paper

OpenStudy (anonymous):

well u hv the value for x,where x=0 and the value of y,where y=-3 and also the slope,m=3 sbstitute this into\[y=mx+c\]to find the value of c

OpenStudy (anonymous):

\[(0,-3)=(x,y)\]\[m=3\]

OpenStudy (anonymous):

\[-3=3(0)+c\]

OpenStudy (anonymous):

what is the value for 3 multiply by 0?

OpenStudy (kweb2017):

3x-y=3

OpenStudy (anonymous):

correct!

OpenStudy (kweb2017):

Thank you :)

OpenStudy (anonymous):

yw ^_^

OpenStudy (anonymous):

Btw,\[\huge{Welcome~To~OpenStudy}\] @kweb2017

OpenStudy (kweb2017):

thank you :)))))

OpenStudy (anonymous):

np ^_^